# Ego-to-Exo: Interfacing Third Person Visuals from Egocentric Views in Real-time for Improved ROV Teleoperation

### v3

@@ Abstract
Underwater ROVs (Remotely Operated Vehicles) are unmanned submersible vehicles designed for exploring and operating in the depths of the ocean. Despite using high-end cameras, typical teleoperation engines based on first-person (egocentric) views limit a surface operator’s ability to maneuver the ROV in complex deep-water missions. In this paper, we present an interactive teleoperation interface that enhances the operational capabilities via increased situational awareness. This is accomplished by (i) offering on-demand “third”-person (exocentric) visuals from past egocentric views, and (ii) facilitating enhanced peripheral information with augmented ROV pose information in real-time. We achieve this by integrating a 3D geometry-based Ego-to-Exo view synthesis algorithm into a monocular SLAM system for accurate trajectory estimation. The proposed closed-form solution only uses past egocentric views from the ROV and a SLAM backbone for pose estimation, which makes it portable to existing ROV platforms. Unlike data-driven solutions, it is invariant to applications and waterbody-specific scenes. We validate the geometric accuracy of the proposed framework through extensive experiments of 2-DOF indoor navigation and 6-DOF underwater cave exploration in challenging low-light conditions. A subjective evaluation on human teleoperators further confirms the effectiveness of the integrated features for improved teleoperation. We demonstrate the benefits of Ego-to-Exo view generation and real-time pose rendering for remote ROV teleoperation by following navigation guides such as cavelines inside underwater caves. This new way of interactive ROV teleoperation opens up promising opportunities for future research in subsea telerobotics.
@@ I Introduction
Unmanned submersible vehicles such as ROVs (Remotely Operated Vehicles) play a crucial role in subsea inspection, remote surveillance, and underwater cave exploration [1, 2, 3]. They are particularly useful in inspecting deep-water structures and surveying confined spaces that are beyond the reach of human scuba divers [4, 5]. In a typical mission, ROVs are controlled by human operators from a surface vessel, who are responsible for the safe and efficient maneuvering of the vehicle [6, 7]. The control consoles for teleoperation typically offer real-time data such as the egocentric video feed, pose, velocity, depth, etc. State-of-the-art ROVs can also include autonomous features for atomic tasks such as hovering [8], following navigation guidelines (e.g., cavelines inside underwater caves [9, 10]), gripping objects [11, 12], trajectory estimation, etc.
@@
While the subsea industries and agencies such as NOAA and naval defense teams deploy underwater ROVs with high-end cameras, sonars, and IMUs [2, 13] – safe and efficient teleoperation remains a challenge in adverse visibility conditions and around complex or sensitive structures. The typical first-person feeds from an ROV camera provide very limited information in landmark-deprived underwater scenes. The operators on the surface can only see the egocentric view, often without global or peripheral semantic information around the ROV. Although ROVs can use artificial lights to enhance visibility in low-light scenes, their bright light get reflected and back-scattered by suspended particles directly at the front camera [9], creating glare and large blind spots for the operator. Additionally, the autonomous and semi-autonomous features of ROVs also become erroneous without peripheral positioning in such noisy sensing conditions.
@@
In this paper, we address these issues by introducing an AR (augmented reality) inspired ROV teleoperation interface that generates third-person (exocentric) perspectives as well as provides interactive control choices for viewpoint selection. As shown in Fig. 1, the proposed console can generate multiple exocentric views from past egocentric images, with a virtual ROV projected on the images as if it were taken by a third person following the robot. Our earlier work introduced the idea of EOB (Eye On the Back) visuals [14], which produces a single third-person view from immediately behind the ROV to facilitate better teleoperation. This work develops the AR-based end-to-end pipeline for generating on-demand exocentric perspectives given any EOB viewpoints. Additionally, we integrate the feature for geometrically accurate ROV positioning into those views for an interactive ROV teleoperation experience in challenging underwater applications. With these features, an operator can slide across multiple EOB viewpoints for the best view, see where the ROV is, and get comprehensive peripheral information for safe and efficient teleoperation.
@@
Specifically, we design an efficient Ego-to-Exo (egocentric to exocentric) view generation framework integrated into a monocular visual SLAM system for underwater ROV teleoperation. The proposed Ego-to-Exo algorithm keeps track of the ROV camera poses and exploits a buffer of egocentric views for exocentric view synthesis. We then transform and project a pre-sampled 3D model of the ROV, in the form of a point cloud, into those views to generate realistic augmented visuals with more peripheral information. As seen in Fig. 1, such views offer comprehensive information of the surrounding scene with global semantics. In addition to the views, the integrated SLAM system provides real-time pose and map updates for atomic tasks [15, 16] such as obstacle avoidance, object following, next-best-view planning, etc.
The significance of this work resides in the simplicity of a 3D-geometric formulation of the Ego-to-Exo problem, and its integration into a real-time SLAM system. A monocular vision-only pipeline is chosen to ensure generalized utility and computational efficiency. The proposed Ego-to-Exo solution, EOB viewpoint-based parameterization, and ROV point cloud projection – are carried out by closed-form solutions to ensure real-time performance. As opposed to data-driven approaches for exocentric view synthesis, the proposed framework is invariant to the changes in waterbody style, scene geometry, and application scenario – making it transferable to any underwater ROV platforms.
We demonstrate the performance and utility of the proposed Ego-to-Exo framework through a series of experiments. First, we prove the geometric validity of our estimation with 2D indoor navigation scenarios utilizing a TurtleBot4. We quantify the geometric validity based on ground plane estimation errors and reprojection errors of known reference points in the scene. Next, we conduct experiments for underwater cave exploration scenarios where a BlueROV2 is teleoperated inside underwater caves and grotto systems at various geographical locations. Remote teleoperation in such cave systems includes unique challenges such as low visibility, turbid water conditions, moving shadow effects, and hazy blind spots due to artificial lights. We evaluate the usefulness of the proposed TeleOp console with human subjects in such challenging scenarios. They use the interactive features such as viewpoint selection to visualize augmented ROV poses on the map in real-time. The subjects rank the complexity, efficiency, and other aspects of the system in a questionnaire as well as provide open-ended feedback regarding their cognitive workload.
We further discuss the observations for a variety of challenging cases of caveline detection and following in noisy low-light conditions inside multiple underwater caves. We demonstrate that the generated exocentric views, the augmented robot pose, and the rendered map embed significantly more information and global context about the scene to facilitate safe and efficient robot control.
II Related Work & Applications
II-A ROV Deployment & Teleoperation
Safe and efficient robot teleoperation is crucial for marine biology, oceanography, offshore energy, and infrastructure maintenance. Underwater ROVs enable researchers to study deep-sea ecosystems and collect samples from depths beyond the limits of human scuba divers. A few prominent applications are as follows.
Subsea structure inspection. Underwater structures operate in challenging remote environments and are susceptible to various hazards with significant economic and environmental consequences. Corrosion/erosion combined with tensile stresses caused by waves and seismic activities lead to cracking, leakage, and malfunction which are regularly inspected by ROVs [1, 17]. ROVs are also deployed for 3D mapping and acoustic profiling of marine structures [18, 19].
Underwater infrastructure maintenance & security. The underwater energy sector [20] and submerged data centers [21, 22] are experiencing rapid growth with increasing demand for regular maintenance and surveillance operations. Once ROVs are teleoperated to the site, they need to perform various autonomous and semi-autonomous tasks to inspect, monitor, and ensure safety and integrity of the infrastructures.
Marine archaeology and cave exploration. Exploring underwater heritage sites [23] and caves [4] are important to unravel archaeological history as well as for water resource management. Intelligent ROV capabilities enable effective mapping, sampling, and exploration of remote sites and challenging overhead environments. Moreover, seasonal monitoring of oceanographic features is performed for photometric studies as well [24, 25, 26].
Beyond these sectors, improved ROV teleoperation interfaces based on our proposed Ego-to-Exo augmented visuals will be useful in coastal habitat restoration and surveying, aquaculture, and search-and-rescue missions as well.
II-B Third-Person Views for ROV Teleoperation
@@
A common issue reported by ROV operators is that using a remote vision platform for teleoperation is like looking through a “soda straw” [27]. This is because the typical ROV controller interfaces are based on egocentric first person camera views – which provide no peripheral vision, resulting in significantly reduced situational awareness [28, 29]. As such, there have been multiple methods proposed to mitigate the issue and provide the operator with more visual data.
@@
Two types of commonly used approaches exist for generating an exocentric view for unmanned ground and aerial vehicles. The first method uses an external camera to provide the exocentric view such as using a fixed camera to record the vehicle’s trajectory from a distance [30], mounting the camera on a UAV (unmanned aerial vehicle) that flies above the target [31, 32, 33, 34], attaching the camera to an elevated position on the robot [35], utilizing a follower ROV with a camera to provide the external view [36], or using a fish-eye camera to create a top-down perspective [37]. The second method utilizes integrated sensors on the robot, such as LiDAR (Light Detection and Ranging) to generate a point cloud of the surrounding environment [38, 39] and use that toward creating an augmented/virtual environment for interfacing and teleoperation [40, 41, 42].
@@
Adapting the aforementioned methods from terrestrial or aerial domains to underwater environments presents inherent challenges. Firstly, sending diver-robot teams [43] is not always an option in complex deep-water missions – which are the majority of use cases for ROVs. Secondly, UGVs that utilize past egocentric views [44, 45] primarily rely on GPS-based localization that does not apply to GPS-denied underwater environment. Unlike underwater ROVs, ground vehicles generally operate on a 2D plane with limited pitch and roll variations over rough terrain. Thirdly, installing an external visual system requires significant hardware modifications, e.g., they need to be rugged and pressure-sealed, re-calibrated for buoyancy and motion dynamics, and additional tether integration for high-speed exocentric data transfer. Even with all the structural modifications, an external camera will provide a single additional third-person perspective. We attempt to address these issues by synthesizing geometrically accurate exocentric views into a monocular SLAM pipeline given a dynamic viewpoint.
III Problem Formulation
@@
We formulate the Ego-to-Exo problem as a 3D geometric algorithm that involves generating an on-demand EOB (Eye On the Back) view, and then projecting the robot on it for an augmented rendering of the scene; see Fig. 2. The proposed method has the following computational components.
III-A Curating Pose and Image Queue from Monocular SLAM
@@
A monocular SLAM algorithm such as ORB-SLAM3 [46] provides a continuous solution for estimating and tracking camera poses from a sequence of monocular images. We use an ORB-SLAM3-based framework to obtain camera poses of each keyframe location to eventually construct the trajectory map of the teleoperated robot. Our adapter SLAM framework initiates the trajectory estimation process by building a pose buffer of length : , where, denotes camera pose at instance in global () frame of reference. The corresponding raw egocentric views for each instance is also stored in a queue . These memory buffers are updated instantaneously as the robot pose changes during teleoperation. We use an empirically tuned threshold to trigger an update only when the pose change is significant to avoid unnecessary updates (when the robot is static).
III-B Generating Exocentric Perspective
@@
Given the pose memory and egocentric views , we formulate the Ego-to-Exo problem of estimating an exocentric view from a reference location , looking toward the robots current location , where and . Typically, is set to (most recent available frame), and remains a free variable with known samples in memory – to mimic the EOB viewpoint generation.
@@
where and represent the pose for current and reference (target) location in world coordinate, respectively. The transformed point cloud is then projected onto the target image plane by using camera intrinsics as:
@@
III-C 3D ROV Rendering and Map Update
While the SLAM system constructs a map of the surrounding, the proposed algorithm simultaneously renders the 3D ROV model on the same spatial context. As illustrated in Fig. 2(b), the ROV points are transformed to the current camera location and projected based on the relative pose information as follows:
@@ IV-A Implementation Details
The framework is implemented using ROS Noetic in an Ubuntu environment, running on an Intel Core i9 processor with GB of RAM. A ROS node for ORB-SLAM3 is integrated as the monocular SLAM backbone. The buffer queue size is set to and the frame separation threshold is set to unit (up to scale). The ROV point clouds are generated by sampling 3D mesh models of BlueROV2 and TurtleBot4; points are sampled for each model. The scaling parameters and are empirically tuned once for each test sequence according to the scale of the SLAM-generated map, ensuring that the rendered ROV model appears realistic in size.

### v5

# EgoExo++: Integrating On-demand Exocentric Visuals with 2.5D Ground Surface Estimation for Interactive Teleoperation of Subsea ROVs

@@ Abstract
Underwater ROVs (Remotely Operated Vehicles) are indispensable for subsea exploration and task execution, yet typical teleoperation engines based on egocentric (first-person) video feeds restrict human operators’ field-of-view and limit precise maneuvering in complex, unstructured underwater environments. To address this, we propose EgoExo, a geometry-driven solution integrated into a visual SLAM pipeline that synthesizes on-demand exocentric (third-person) views from egocentric camera feeds. Our proposed framework, EgoExo++, extends beyond 2D exocentric view synthesis (EgoExo) to augment a dense 2.5D ground surface estimation on-the-fly. It simultaneously renders the ROV model onto this reconstructed surface, enhancing semantic perception and depth comprehension. The computations involved are closed-form and rely solely on egocentric views and monocular SLAM estimates, which makes it portable across existing teleoperation engines and robust to varying waterbody characteristics. We validate the geometric accuracy of our approach through extensive experiments of 2-DOF indoor navigation and 6-DOF underwater cave exploration in challenging low-light conditions. Quantitative metrics confirm the reliability of the rendered Exo views, while a user study involving operators demonstrates improved situational awareness, navigation safety, and task efficiency during teleoperation. Furthermore, we highlight the role of EgoExo++ augmented visuals in supporting shared autonomy, operator training, and embodied teleoperation. This new interactive approach to ROV teleoperation presents promising opportunities for future research in subsea telerobotics.
@@ I Introduction
Unmanned submersible vehicles such as ROVs (Remotely Operated Vehicles) play a crucial role in subsea inspection, remote surveillance, and underwater cave exploration [1, 2, 3]. They are particularly useful in inspecting deep-water structures and surveying confined spaces that are beyond the reach of human scuba divers [4, 5]. In a typical mission, ROVs are controlled by human operators from a surface vessel, who are responsible for the safe and efficient maneuvering of the vehicle [6, 7]. The control consoles for teleoperation typically offer real-time data such as the egocentric video feed, pose, velocity, depth, etc. State-of-the-art (SOTA) ROVs can also include autonomous features for atomic tasks such as hovering [8], following navigation guidelines inside underwater caves and overhead structures [9, 10, 11], object manipulation [12, 13], trajectory estimation, etc.
@@
While the subsea industries and agencies such as NOAA and naval defense teams deploy underwater ROVs with high-end cameras, sonars, and IMUs [2, 14] – safe and efficient teleoperation remains a challenge in adverse visibility conditions and around complex or sensitive structures. The typical first-person feeds from an ROV camera provide very limited information in landmark-deprived underwater scenes. The operators on the surface can only see the egocentric view, often without global or peripheral semantic information around the ROV [15, 16]. Although ROVs can use artificial lights to enhance visibility in low-light scenes, their bright light get reflected and back-scattered by suspended particles directly at the front camera, creating glare and large blind spots for the operator [10]. Additionally, the autonomous and semi-autonomous features of ROVs become erroneous without peripheral positioning in such noisy sensing conditions.
@@
In this paper, we address these issues by introducing an AR (augmented reality) inspired ROV teleoperation interface that generates third-person (exocentric) perspectives as well as provides interactive control choices for viewpoint selection. As shown in Figure 1, the proposed console can generate multiple exocentric views from past egocentric images, with a virtual ROV model projected on the images as if it were taken by a third person following the robot. Our early work introduced the idea of EOB (Eye On the Back) visuals [18], envisioning a single third-person view from immediately behind the ROV to facilitate better teleoperation. Our recent work materialized this idea in EgoExo [17], by formalizing an AR-based framework that generates on-demand exocentric imagery from any EOB viewpoint. It also integrated the feature for geometrically accurate ROV positioning into those views. This work further advances this direction of research by introducing EgoExo++, a dynamic 2.5D exocentric visualization – analogous to a bird’s-eye view in terrestrial contexts – that offers an interactive and semantically enriched perspective of the environment. Importantly, our approach is closed-form and solely geometry-driven, ensuring accuracy and real-time efficiency without reliance on data-driven methods or training biases. The envisioned interface supports both fore-aft transitions across multiple EOB views and an interactive, rotatable exocentric perspective, enabling a safer and more informed ROV teleoperation.
@@
Specifically, we introduce an efficient framework for generating egocentric to exocentric (and beyond) visual perspectives integrated into a visual SLAM system for underwater ROV teleoperation. The base EgoExo algorithm keeps track of the ROV camera poses, visual features extracted by SLAM, and exploits a buffer of egocentric views for exocentric view synthesis. We then transform and project a pre-sampled 3D model of the ROV, in the form of a point cloud, into those views to generate realistic augmented visuals with more peripheral information. In parallel, the EgoExo++ pipeline utilizes SLAM-generated feature points to identify ground regions and fuses them into a continuous ground surface where pixel colors are transferred from corresponding image regions. We employ a temporal fuse-and-stack strategy to preserve the historical ground evidence, while the 3D ROV model is projected on the same spatial context. The resulting 2.5D perspective enables operators to interact with the scene using dynamic viewpoints in real-time. As illustrated in Figure 1, these views provide operators with a globally informed and semantically rich understanding of the surrounding environment, supporting interactive control from arbitrary viewpoints. In addition to view synthesis, the SLAM backend delivers real-time updates on camera pose and environmental mapping to better assist with atomic tasks [13, 19, 20] such as obstacle avoidance, object following, next-best-view planning, manipulation, etc.
II Background and Related Work
II-A Third-Person Views for ROV Teleoperation
@@
A common issue reported by ROV operators is that using a remote vision platform for teleoperation is like looking through a “soda straw” [21, 18]. This is because the typical ROV controller interfaces are based on egocentric first person camera views – which provide no peripheral vision, resulting in significantly reduced situational awareness [22, 23]. Researchers have explored both fixed [24, 25] and dynamic [26, 27] viewpoint augmentation methods in contemporary human-machine interface study [28, 29].
@@
Two primary approaches are used for generating exocentric views in unmanned ground and aerial vehicles. The first leverages external cameras to capture the vehicle’s motion from a distance; examples include fixed ground cameras [30], UAV-mounted overhead views [31, 32, 33, 34], elevated on-robot mounts [35], camera-equipped follower ROVs [36], and fisheye lenses for top-down perspectives [37, 38]. The second method utilizes additional onboard sensors, such as LiDAR (Light Detection and Ranging), to generate a point cloud of the surrounding environment [24, 25] and use it to create an augmented/virtual reality for interfacing and teleoperation [39, 38, 40, 41].
@@
Adapting the aforementioned methods from terrestrial or aerial domains to underwater environments presents inherent challenges. Firstly, sending diver-robot teams [42] is not always an option in complex deep-water missions – which are the majority of use cases for ROVs. Secondly, UGVs that utilize past egocentric views [43, 44] primarily rely on GPS-based localization that does not apply to GPS-denied underwater environments. Unlike underwater ROVs, ground vehicles generally operate on a 2D plane with limited pitch and roll variations over rough terrain [45]. Thirdly, installing an external visual system requires significant hardware modifications, e.g., they need to be rugged and pressure-sealed, recalibrated for buoyancy and motion dynamics, and additional tether integration for high-speed data transfer. Even with all the structural modifications, an external camera will provide a single additional third-person perspective.
A range of AR/VR-based teleoperation systems have been developed to enhance operator immersion and augment visual feedback for subsea tasks such as object grasping and manipulation [46, 29, 13, 47], inspection [48, 49], and navigation [50]. These systems commonly support third-person perspectives by embedding the operator within an XR (extended reality) environment that incorporates a digital twin of the ROV [51] and reconstructs the surrounding scene using 3D models. While such immersive interfaces improve situational awareness and control, they often require extensive sensory augmentation (e.g., visual, auditory, haptic) at both the ROV and operator ends [13, 41], which increases hardware demands and complicates real-time deployment.
II-B 2.5D Exocentric View Generation
Generating 2.5D/3D third-person views from front-facing camera is critical for scene understanding, both for human teleoperators and for autonomous vehicles. The challenges lie in extreme viewpoint shift and lack of direct depth cues from monocular inputs. Recent efforts for 2.5D view synthesis can be categorized into two main areas: homography-based geometric projections [52, 53] and generative models using encoder-decoder, adversarial, or transformer-based learning [54, 55].
The geometry-guided CNN proposed by [53] warps frontal images to the top view using a fitted homography matrix. While efficient for structured environments, their approach is limited to the flat-ground assumption and struggles with non-planar surfaces. Zhu et al. [56] introduce an intermediate homography view from generative adversarial network (GAN) to reduce the difficulty of pure geometric transformation. Transformer models such as BEVFormer [54] and BEVDepth [57] integrate temporal or multi-view cues to improve realism in synthesized views at a cost of high computation. Other learning-based approaches fuse multiple camera views or additional sensors (e.g., LiDAR) to generate semantic aerial views [58, 59], diverging from monocular egocentric setups. Unlike these data-driven approaches, we propose a lightweight geometric solution, integrated into a visual SLAM pipeline that offers real-time, interactive third-person perspectives without relying on multi-modal sensory augmentation or additional hardware.
III EgoExo++: Problem Formulation
@@
We formulate the EgoExo++ problem as a 3D geometric algorithm that involves generating an on-demand EOB view, reconstructing the ground surface, and then projecting the ROV model both on 2D and 2.5D context for augmented rendering of the scene; see Figure 2. The proposed method has the following computational components.
III-A Curating ROV Pose and Image Buffer
@@
A monocular SLAM algorithm such as ORB-SLAM3 [60] provides a continuous solution for estimating and tracking camera poses from a sequence of monocular images. We use an ORB-SLAM3-based framework to obtain camera poses of each keyframe location to eventually construct the trajectory map of the teleoperated robot. In our implementation, the SLAM pipeline initiates the trajectory estimation process by building a pose buffer of length : , where, denotes camera pose at instance in global () frame of reference. The corresponding raw egocentric views for each instance is also stored in a queue . These memory buffers are updated instantaneously as the robot pose changes during teleoperation. We use an empirically tuned threshold to trigger an update only when the pose change is significant to avoid unnecessary updates (when the robot is static).
III-B Generating 2D Exo Image
@@
Given the pose memory and egocentric views , we formulate the EgoExo problem of estimating an exocentric view from a reference location , looking toward the robot’s current location , where and . Typically, is set to (most recent available frame), and remains a free variable with known samples in memory – to mimic the EOB viewpoint generation.
@@
where and represent the ROV pose for current and reference (target) location in world coordinate, respectively. The transformed point cloud is then projected onto the target image plane by using camera intrinsics as:
@@
III-C Generating 2.5D Exo Views
In EgoExo++, we reuse the SLAM-generated visual features from the EgoExo pipeline to estimate the ground surface and synthesize a lightweight terrain-aware 2.5D perspective. This process involves four stages: (i) selecting candidate feature points for the ground surface, (ii) fitting the ground plane, (iii) translating texture from image pixels to the estimated surface, and (iv) fusing multiple frames over time for real-time visualization.
Due to the lack of horizon line in open water settings and the uneven geometry of confined underwater spaces (e.g., caves), we incorporate geometric priors based on the camera orientation to initialize the ground region estimation. In the nominal case with zero pitch and roll, the ground remains within the bottom half of the image, separated by a horizontal line at (where is the image height). As the camera pitches downward, this line shifts upward, since a larger portion of the ground comes within the camera’s FOV, and vice versa. A camera roll rotates this dividing line accordingly. By computing this orientation-adjusted imaginary horizon from the known camera pose, we restrict candidates to points that fall within the “ground side” of the image. This prior ensures that no 3D point projecting above the horizon (e.g., from cave walls and ceiling) is selected as ground.
Let be the set of SLAM feature points in the world frame, associated with camera pose at time instance . After imposing the geometric prior and pre-selecting candidate points, we fit a plane via RANSAC [61]:
where is an inlier loss with threshold . To enforce stability, we apply a prior that constrains the plane normal within an angle of the expected vertical direction ( in camera frame). Given the plane and a reference anchor (closest point from camera center to ), we define an orthonormal basis on the plane. Each 3D point is expressed in local coordinates as:
A rectangular grid is constructed on the ground plane, and the sparse heights are interpolated to obtain a smooth elevation field . Each grid vertex
is then reprojected to the image using intrinsics :
Image colors are sampled (bilinear interpolation) to texture the grid, producing a dense D ground surface. To extend the ground beyond a single camera frame, all historical ground patches are accumulated in the global frame. Patches are merged using voxel decimation and Delaunay triangulation [62] to avoid redundancy while preserving continuity. The fused mesh forms a D exocentric perspective with surface geometry and realistic coloring consistent with the egocentric imagery.
III-D ROV Model Rendering and Scene Update
While the SLAM system constructs a sparse map of the surroundings, the proposed algorithm simultaneously renders the 3D ROV point cloud (or mesh for EgoExo++) on the same spatial context. The ROV points are transformed to the current camera location and projected based on the relative pose information as follows:
@@ IV-A Implementation Details
The framework is implemented using ROS Noetic in an Ubuntu environment, running on an Intel Core i9 processor with GB of RAM. A ROS node for ORB-SLAM3 is integrated as the monocular SLAM backbone. The buffer queue size is set to and the frame separation threshold is set to unit (up to scale). The ROV point clouds are generated by sampling 3D mesh models of BlueROV2 and TurtleBot4; points are sampled for each model. The scaling parameters , , and are empirically tuned once for each test sequence according to the scale of the SLAM-generated map, ensuring that the rendered ROV model appears realistic in size. Note that we adopt the North-East-Down (NED) frame convention used by [65], which is local to the SLAM origin (not aligned with Earth’s North/East).